OpenStudy (anonymous):

Solve the nonlinear inequality 3+x/3-x greater or equal 1

OpenStudy (anonymous):

Help

OpenStudy (anonymous):

\[\frac{ 3+x }{ 3-x }\ge1\] multiply both sides by this \[(3-x)^2\]

OpenStudy (anonymous):

Why is it squared

OpenStudy (anonymous):

Because you don't know what x is.

OpenStudy (anonymous):

its not, I think he typed that by accident

OpenStudy (anonymous):

it could be 9 and then you just ,mulktiplyiing both sides by a negative.

OpenStudy (anonymous):

multiplying.

OpenStudy (anonymous):

no....@sweet1137

OpenStudy (anonymous):

Kk

OpenStudy (anonymous):

no no that's not correct

OpenStudy (anonymous):

I'm confused

OpenStudy (anonymous):

Here's an example.

OpenStudy (anonymous):

hold on, both of you

OpenStudy (anonymous):

if x is 9 in your inequality.

OpenStudy (anonymous):

\[\frac{ 3+9 }{ 3-9 }\ge 1\] And you multiply by 3-9 to both sides.

OpenStudy (anonymous):

\[\frac{ 12 }{ -6 }\times -6 \ge 1\times -6\]

OpenStudy (anonymous):

you must change the inequalitysign.

OpenStudy (anonymous):

that solves the inequality, x cannot be 3, since for x=3 the expression is undefined yet for all other values of x greater than 0, the expression will yield values larger than 1

OpenStudy (anonymous):

you're right

OpenStudy (anonymous):

no...you must multiply by the square of the denominator so the inequality sign won't get affected...

OpenStudy (anonymous):

my b, my b

OpenStudy (anonymous):

yes, you're right

OpenStudy (anonymous):

So what do I do??.

OpenStudy (anonymous):

\[\frac{ (x+3)(x-3)^2 }{ x-3 }\ge (x-3)^2\] \[(x+3)(x-3)\ge (x-3)^2\] \[(x+3)(x-3)-(x-3)^2\ge 0\] \[(x-3)[x+3-(x-3)]\ge 0\] \[6(x-3)\ge 0\]

OpenStudy (anonymous):

that still doesn't work, since x=0 is a solution not included in that inequality

OpenStudy (anonymous):

Jeez

OpenStudy (anonymous):

\[x-3\ge 0\] 

OpenStudy (anonymous):

SInce the solution would be x>3 since x cannot be 3

OpenStudy (anonymous):

If x needs to be greater than 3, like your inequality states, why is it that for x = 0, 3+0/3-0 is greater than or equal to 1

OpenStudy (anonymous):

But then you're right. I'm trying think back to my early days of doing this stuff.

OpenStudy (anonymous):

hahah yea man, I'm trying too

OpenStudy (anonymous):

Made a ,mistake

OpenStudy (anonymous):

\[(3-x)(3+x-3+x)\ge 0\]

OpenStudy (anonymous):

two solutions. \[2x(3-x)\ge 0\]

OpenStudy (anonymous):

Solved.

OpenStudy (anonymous):

I did a boo boo.

OpenStudy (anonymous):

How did u do that lmao

OpenStudy (anonymous):